OpenStudy (blaque678):

Find all solutions, real and complex of the equation x^3=1

OpenStudy (tkhunny):

This is not all that tricky once you have seen it. Such complex roots are spaced evenly around the origin. Since we seek three, we'll need \(\pi/3\) or 120º between each one. Since also there is a trivial solution, x = 1, the other two are not hard to find.

OpenStudy (anonymous):

\[\frac{2\pi}{3} \]

OpenStudy (luigi0210):

I hope you're not giving an answer pg

OpenStudy (anonymous):

uh... no just fixing a boo boo

OpenStudy (luigi0210):

Just making sure :P

OpenStudy (tkhunny):

Let's see, \(\2\pi / 3 = \dfrac{2\pi}{3}\). What do you know! Sorry about that. Boo boo acknowledged and repaired.

OpenStudy (tkhunny):

Okay, my typing just keeps getting worse. Might be time for a break.

OpenStudy (anonymous):

no worries!

OpenStudy (blaque678):

I don't know how to the problem at all

OpenStudy (tkhunny):

There are at least two ways to do this one. One pretty fancy. I was trying to hint at it, above. We also have algebra. To find the solutions to x^3 = 1, we solve x^3 - 1 = 0. Are we feeling better about that "at all" thing? It is hoped that you have factored and solved such equations, before.

OpenStudy (blaque678):

ok so x=1 and the other two are a quadratic equations?

OpenStudy (tkhunny):

Yes, you can use the quadratic formula to find the other two. You can think to the future at the exciting other ways there are to find such solutions. Good work. I invite you to plot these three solutions and ponder on the first answer. See how they are evenly spaced around a circle?